# BConformeR: A Conformer Based on Mutual Sampling for Unified Prediction of Continuous and Discontinuous Antibody Binding Sites

### v2

@@ 4.1.1 Dataset
Table 2: Performance comparison of different methods on a blind test of 24 antigen records. F1-L and F1-D refer to the F1 scores calculated on linear and discontinuous epitopes, respectively; Thresholds were selected following those of baseline methods (Appendix E.2); Standard deviation calculated over a range of thresholds.
Due to the differences in training and test datasets across existing epitope prediction methods, the selection of a blind test set was important for fair evaluation. Therefore, we also extracted 24 antigen records from the test sets of SEMA-1D 2.0 and DiscoTope-3.0, all of which are blind to BConformeR.
Model Linear Discontinuous Params (M) MACs (G)
Precision Recall F1 Precision Recall F1
BConformer-12 0.440 0.477 0.458 0.304 0.360 0.330 302.24 309.93
BConformer-9 0.428 0.463 0.445(-2.8%) 0.297 0.322 0.309(-6.4%) 230.27 236.12
BConformer-6 0.423 0.416 0.419(-8.5%) 0.249 0.296 0.270(-18.2%) 158.30 162.31
ResNet1D-152 0.442 0.484 0.462(+0.9%) 0.255 0.291 0.272(-17.6%) 242.44 248.62
ResNet1D-101 0.417 0.430 0.423(-7.6%) 0.242 0.238 0.240(-27.3%) 179.24 183.80
ViT1D-12 0.429 0.435 0.432(-5.7%) 0.269 0.334 0.298(-9.7%) 230.24 234.42
ViT1D-9 0.403 0.428 0.415(-9.4%) 0.276 0.311 0.292(-11.5%) 173.57 176.32
Table 3: Ablation study of different models evaluated on 100 antigen records (Appendix E.1). BConformer-12 corresponds to BConformeR, and BConformer-9/6 reduces one layer of FCU-based CNN and Transformer for each stage compared to BConformer-12/9. ResNet bottlenecks adopt the same CNN blocks as used in Conformer, and ViT follows the design of Transformer branch (Appendix F). Models are grouped with comparable Params and MACs: BConformer-9, ResNet1D-152 and ViT1D-12; BConformer-6, ResNet1D-101 and ViT1D-9.
4.1.2 Data encodings and features
To represent antigen sequences, we used pretrained ESM-2 embeddings. Sequences were tokenized with the ESM-2 tokenizer and passed through its 33-layer transformer. We extracted the hidden states from the final layer, discarded the beginning-of-sequence (BOS) and end-of-sequence (EOS) tokens, and retained only the embeddings corresponding to the actual residues. Finally, each antigen sequence was represented as a $\mathbb{R}^{1024\times 1280}$ feature matrix, with 1024 as the padded length and 1280 as the ESM-2 embedding size.
@@ 4.2.1 Benchmark evaluation
The performance of BConformeR was evaluated on a blind test set of 24 antigen records and compared with four SOTA methods: SEPPA-3.0, BepiPred-3.0, SEMA-1D 2.0 and DiscoTope-3.0. As shown in Table 2, BConformeR outperforms baselines in MCC (0.296), PCC (0.347), and AgIoU (0.203). Further evaluation of ROC-AUC and PR-AUC shows that BConformeR achieves scores of 0.777 and 0.313 respectively, surpassing all methods that provided residue-level probabilistic outputs (Fig. 3).
4.2.2 Enhanced prediction on continuous and discontinuous epitopes
To evaluate the ability of each method in predicting continuous and discontinuous epitopes, we computed the F1 scores separately on the test set. As shown in Table 2, BConformeR achieves F1 scores of 0.349 and 0.262 on linear and discontinuous epitopes, respectively, higher than all baseline methods. This illustrates that BConformeR is more effective at identifying both regions and fragments of epitopes.
@@ 4.3.1 Overview
4.3.2 Depth reduction
To evaluate the impact of model depth, we progressively reduced the number of layers in the Conformer architecture. As shown in Table 3, reducing depth to BConformer-9 leads to a moderate drop in performance, with linear and discontinuous F1 scores decreasing by 2.8% (0.445) and 6.4% (0.309) respectively. Further reduction to BConformer-6 causes a more significant decline, especially in discontinuous epitope prediction, where the F1 score drops by 18.2% to 0.270. These demonstrate that our final model, BConformeR, achieves superior performance compared to shallower variants.
@@ 4.3.2 Depth reduction
Purely convolutional variants capture contiguous residue patterns efficiently and therefore achieve competitive scores on linear epitopes. However, their strictly local receptive field prevents them from modelling long-range residue couplings that dominate conformational epitopes. As shown in Table 3, ResNet1D-152 attains the highest linear F1 score (0.462) among all models, confirming its effectiveness on local patterns. However, its discontinuous F1 drops to 0.272, lower than BConformer-9 (0.309) and ViT1D-12 (0.298), pointing to difficulty in capturing discontinuous dependencies.
@@
4.4 High-Confidence Predicted Structures
@@ 4.4.1 Predicted epitopes with high AgIoU
5 Conclusion and Limitation
We present BConformeR, a dual–branch Conformer model that integrates convolutional and self–attention mechanisms for residue–level prediction of both linear and conformational B-cell epitopes from antigen sequences. In addition to the architecture itself, two auxiliary contributions are worth noting: (i) an adaptive logit–fusion mechanism that automatically balances evidence from the CNN and Transformer branches, yielding a single, well-calibrated probability for each residue without extra post-processing; (ii) a detailed performance analysis on both linear and discontinuous epitopes, which also validates the complementary effectiveness of the CNN and Transformer branches. Benchmark evaluation demonstrates that our model consistently outperforms strong sequence- and structure-based baselines, particularly on the more difficult task of predicting conformational epitopes. The anticipated beneficiaries include rational vaccine design, diagnostic reagent development and antibody engineering pipelines.
Limitations.
There are three notable limitations to this study. First, all training and test interfaces are derived from in silico AlphaFold-Multimer models; although this choice enlarges the dataset, it may introduce misoriented chains or misplaced atoms. Second, antigen families and epitope types remain imbalanced, which could bias both training and evaluation. This is reflected in the significantly higher F1 scores observed on the 100-antigen test set compared with the 24-antigen blind test set, even though all records are unseen by BConformeR. Third, considering the computational cost, input sequences are limited to 1024 residues. Antigens longer than this threshold (e.g., 2nyy, 3ri5) cannot be directly processed by the model.
@@ Acknowledgments
Figure 5: Related work on epitope prediction methods. FFNN denotes feedforward neural network; FC layers denotes fully connected layers; MSA denotes multiple sequence alignment results.
@@
Figure 6: Linear regions and discontinuous fragments. (A) (Left) The blue structure represents antigen 7c88, and the light green and light orange correspond to part of the antibody light chain and heavy chain, respectively; Linear regions are marked with red rectangles, and discontinuous fragments with green. (Right) Residues highlighted with numbers indicate epitopes. (B) A linear region of 7c88. Gap size of up to three non-contacting residues is chosen for a linear region on the basis of the structure of $\alpha$-helices.
Appendix A Antibody-agnostic prediction methods
We summarize widely-used epitope prediction methods, categorized into sequence-based and structure-based approaches, as shown in Figure 5. Methods with the latest versions are summarized along with their data representation strategies and classifiers.
Appendix B Linear Regions and Discontinuous Fragments
@@
Regions are defined as linear stretches of antigen sequence having at least three antibody-contacting residues (Fig. 6A). Gaps between contacting residues are allowed, and a gap size of up to three non-contacting residues is chosen on the basis of the structure of $\alpha$-helices, which enables residues on the same helical face to be included (Fig. 6B). Fragments, which do not contain enough antibody-contacting residues to qualify as a region (less than three), are considered discontinuous in our study (Fig. 6A). An algorithm classifying linear epitopes for an antigen sequence is shown below:
Algorithm 1 Linear Epitope Classification
1:label_tensorlabel\_tensor : sequence of residue labels
2:   ($1=$ contacting, $0=$ non-contacting, 100=-100= padding)
3:linear_flagslinear\_flags: boolean mask indicating linear epitope residues
4:Convert label_tensorlabel\_tensor to list labelslabels of length $L$
5:Initialize $linear\_flags\leftarrow[\text{False}]^{L}$, $visited\leftarrow[\text{False}]^{L}$
[…]
7:while $i<L$ do
8:  if $labels[i]\neq 1$ or $visited[i]=$ True then
9:   $i\leftarrow i+1$ $\triangleright$ skip non-epitope or already visited residue
10:   continue
11:  end if
12:  Initialize $segment\_indices\leftarrow[\,]$, $one\_count\leftarrow 0$, $zero\_count\leftarrow 0$
13:  $j\leftarrow i$
14:  while $j<L$ do
15:   if labels[j]=100labels[j]=-100 then
16:     break $\triangleright$ stop at padding
17:   else if $labels[j]=1$ then
18:     Append $j$ to segment_indicessegment\_indices
19:     $one\_count\leftarrow one\_count+1$, $zero\_count\leftarrow 0$
20:   else if $labels[j]=0$ then
21:     $zero\_count\leftarrow zero\_count+1$
22:     if zero_count>3zero\_count>3 then
23:      break $\triangleright$ gap too large
24:     end if
25:     Append $j$ to segment_indicessegment\_indices
26:   end if
27:   $j\leftarrow j+1$
28:  end while
29:  $ones\_in\_segment\leftarrow\{k\in segment\_indices\mid labels[k]=1\}$
30:  if $|ones\_in\_segment|\geq 3$ then
31:   for $k\in ones\_in\_segment$ do
32:     $linear\_flags[k]\leftarrow\text{True}$, $visited[k]\leftarrow\text{True}$
33:   end for
34:   $i\leftarrow segment\_indices[-1]+1$
35:  else
36:   $visited[i]\leftarrow\text{True}$, $i\leftarrow i+1$
37:  end if
38:end while
39:return linear_flagslinear\_flags
Appendix C Dataset
C.1 Structure prediction
We initially obtained 1730 Ag-Ab FASTA files from AbEpiTope-1.0. However, due to the limited computational time and the large size of certain input sequences, we only processed 1300 complexes using the AlphaFold2 Multimer v3. Each of these 1300 complexes was modeled with 5 seeds, 6 models and 3 recycles, generating 30 structures per complex.
C.2 Epitope statistics
The final dataset consists of 1180 Ag–Ab complexes, which contains a total of 1432 antigen chains and 2360 antibody chains. Epitopes were identified using a 4 Å distance cutoff between antigen and antibody residues. After classifying the epitopes into linear and discontinuous types, we found 2388 linear regions comprising 15,886 residues, and 2632 discontinuous fragments comprising 3574 residues.
Appendix D Performance Metrics
@@
The performance metrics used in this study include Precision, Recall, True Positive Rate (TPR), False Positive Rate (FPR), F1-score, Antigen-level intersection over union (AgIoU), Matthews correlation coefficient (MCC) and Pearson correlation coefficient (PCC), whose definitions are given as follows:
$\text{Precision}=\frac{TP}{TP+FP},$
$\text{Recall}=\frac{TP}{TP+FN},$
$\text{TPR}=\frac{TP}{TP+FN},$
$\text{FPR}=\frac{FP}{FP+TN},$
$\text{F1}=\frac{2TP}{2TP+FP+FN},$
$\text{AgIoU}=\frac{TP}{TP+FP+FN},$
$\text{MCC}=\frac{TP\cdot TN-FP\cdot FN}{\sqrt{(TP+FP)(TP+FN)(TN+FP)(TN+FN)}},$
$\text{PCC}=\frac{\sum_{i=1}^{n}(y_{i}-\bar{y})(\hat{y}_{i}-\overline{\hat{y}})}{\sqrt{\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}}\sqrt{\sum_{i=1}^{n}(\hat{y}_{i}-\overline{\hat{y}})^{2}}},$
@@
where $TP$, TNTN, FPFP and $FN$ denote the number of true positives, true negatives, false positives and false negatives, respectively; $y_{i}$ and $\hat{y}_{i}$ represent the ground-truth and predicted values, while $\bar{y}$ and $\overline{\hat{y}}$ denote their means. Note that Recall and TPR are mathematically equivalent, but used in different contexts. The relationships between some of these metrics are given as:
$\text{F1}=\frac{2\cdot\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}=\frac{2\cdot\text{AgIoU}}{1+\text{AgIoU}},$
@@
Appendix E Evaluation and Comparison
E.1 Evaluation
Figure 7: Metrics across thresholds on the test set of 100 antigens.
BConformeR was evaluated at various thresholds on 100 antigen records. As shown in Figure 7, the range between 0.1 and 0.5—highlighted by gray dashed lines—represents the region where precision, recall, F1 score and AgIoU perform well. Notably, the F1 score peaks around 0.44 when thresholds are set between 0.1 and 0.3.
E.2 Baseline methods
Thresholds for all baselines were adjusted to reflect their best scoring behavior, ensuring a fair comparison with BConformeR in the blind test. We observed that SEPPA-3.0 performs better under a low decision threshold (around 0.1). For SEMA-1D 2.0, which defines epitope residues using an 8 Å distance cutoff, the predicted scores tend to be higher; thus, we adopted a threshold of 0.5 above its default of 0.351. Moreover, although the default threshold of BepiPred-3.0 is 0.1512, we found that a slightly higher value around 0.22 led to better balance between precision and recall on our test set.
Appendix F Ablation
F.1 ViT1D
Model Embed Dimension Seq Length Patch Size Depth Heads MLP Ratio
ViT1D-9 1536 1024 1 9 12 2.0
ViT1D-12 1536 1024 1 12 12 2.0
@@ F.2 ResNet1D
input 1280 - $\checkmark$ $\checkmark$
stem 320 7 $\checkmark$ $\checkmark$
layer1 640 (160 $\times$ 4) 3 3 3
layer2 1280 (320 $\times$ 4) 3 4 8
layer3 2560 (640 $\times$ 4) 3 23 36
layer4 5120 (1280 $\times$ 4) 3 3 3
@@
We employed two variants of the ResNet1D backbone in the ablation studies: ResNet1D-101 and ResNet1D-152. As shown in Table 5, both models accept input feature maps with 1280 channels and begin with an initial convolutional layer producing 320 channels. The networks consist of four residual stages of bottlenecks with the expansion factor of 4, each stage starting by downsampling the channel dimension to half. The difference lies in the number of bottleneck blocks per stage. ResNet1D-101 contains [3, 4, 23, 3] blocks across the four layers, while ResNet1D-152 is deeper with [3, 8, 36, 3] blocks. The final output is projected to the target classes through a $\times1$ convolutional layer.

### v3

@@ 3.1.2 Calibrated score
Table 2: Performance comparison of different methods. F1-L and F1-D refer to the F1 scores calculated on linear and discontinuous epitopes, respectively. Standard deviation calculated over a range of thresholds.
@@ 3.2.1 Benchmark evaluation
BConformeR outperforms state-of-the-art baselines, achieving superior ROC-AUC and PR-AUC on Alphafold-predicted structures (Fig. 3E,F) and higher PR-AUC on SARS-CoV-2 test set (Fig. 3G). Notably, it consistently improves the prediction of both linear and discontinuous epitopes (Tab. 2), demonstrating robust identification of epitope regions and fragments.
Model Linear Discontinuous Params (M) MACs (G)
Prec $\uparrow$ Rec $\uparrow$ F1 $\uparrow$ Prec $\uparrow$ Rec $\uparrow$ F1 $\uparrow$
BConformer-12 0.440 0.477 0.458 0.304 0.360 0.330 289.31 296.64
BConformer-9 0.428 0.463 0.445(-2.8%) 0.297 0.322 0.309(-6.4%) 219.94 225.50
BConformer-6 0.423 0.416 0.419(-8.5%) 0.249 0.296 0.270(-18.2%) 150.56 154.37
ResNet1D-152 0.442 0.484 0.462(+0.9%) 0.255 0.291 0.272(-17.6%) 196.41 201.45
ResNet1D-101 0.417 0.430 0.423(-7.6%) 0.242 0.238 0.240(-27.3%) 145.21 148.92
ViT1D-12 0.429 0.435 0.432(-5.7%) 0.269 0.334 0.298(-9.7%) 230.05 234.22
ViT1D-9 0.403 0.428 0.415(-9.4%) 0.276 0.311 0.292(-11.5%) 173.37 176.12
Table 3: Ablation studies evaluated on the validation set. BConformer-12 corresponds to BConformeR, and BConformer-9/6 reduces one layer of FCU-based CNN and Transformer for each stage compared to BConformer-12/9. ResNet bottlenecks adopt the same CNN blocks as Conformer, and ViT follows the design of Transformer branch.
@@ 3.3.1 Overview
Purely convolutional variants capture contiguous residue patterns efficiently and therefore achieve competitive scores on linear epitopes. However, their strictly local receptive field prevents them from modelling long‐range residue couplings that dominate conformational epitopes. As shown in Table 3, ResNet1D-152 attains the highest linear F1 score (0.462) among all models, confirming its effectiveness on local patterns. However, its discontinuous F1 drops to 0.272, lower than BConformer-9 (0.309) and ViT1D-12 (0.298), pointing to difficulty in capturing discontinuous dependencies.
@@
3.4 Case Studies
@@ 3.4.1 Predicted epitopes with high AgIoU
4 Related Work
Numerous antibody-agnostic epitope prediction methods have been proposed. SEPPA-3.0 was the epitope prediction tool specifically designed for glycoproteins, based on the logistic regression classifier SEPPA3.0. BepiPred-3.0 used embeddings from pretrained protein language models ESM-2 as input to a feedforward neural network (FFNN) to predict both linear and conformational B-cell epitopes BepiPred3.0. Epitope3D modeled antigen surfaces as residue-level graphs and extracted distance-based graph signatures to train classifiers for conformational B-cell epitope prediction, with AdaBoost selected as the backbone Epitope3D; AdaBoost. ScanNet was an end-to-end geometric deep learning model that learned interpretable residue-level features directly from 3D structural neighborhoods ScanNet. SEMA-1D 2.0 was constructed by adding a fully-connected linear layer on an ensemble of five ESM-2 models, while SEMA-3D 2.0 was developed upon an ensemble of five pre-trained SaProt models SEMA2.0; SaProt. DiscoTope-3.0 utilized inverse folding representations from ESM-IF1 and was trained on both predicted and solved structures using a positive-unlabelled ensemble strategy, enabling structure-based B-cell epitope prediction at scale DiscoTope3.0; ESM-IF1.
5 Conclusion
We present BConformeR, a dual–branch Conformer model that integrates convolutional and self–attention mechanisms for residue–level prediction of both linear and conformational B-cell epitopes from antigen sequences. Two contributions are worth noting: (i) Adaptive feature coupling and logit fusion that balance information from the CNN and Transformer branches, yielding a well-calibrated score for each residue; (ii) a detailed performance analysis on both linear and discontinuous epitopes, which also validates the complementary effectiveness of both branches. Benchmark evaluation demonstrates that our model consistently outperforms strong sequence- and structure-based baselines, particularly on the more difficult task of predicting conformational epitopes. The anticipated beneficiaries include rational vaccine design, diagnostic reagent development and antibody engineering pipelines.
@@ Acknowledgments
Figure 5: Linear regions and discontinuous fragments. (Left) The blue structure represents antigen 7c88, and the light green and light orange correspond to part of the antibody light chain and heavy chain, respectively; Linear regions are marked with red rectangles, and discontinuous fragments with green. (Right) Residues highlighted with numbers indicate epitopes.
Appendix A Linear Regions and Discontinuous Fragments
@@
Regions are defined as linear stretches of antigen sequence having at least three antibody-contacting residues. Gaps between contacting residues are allowed, and a gap size of up to three non-contacting residues is chosen on the basis of the structure of $\alpha$-helices, which enables residues on the same helical face to be included. Fragments, which do not contain enough antibody-contacting residues to qualify as a region (less than three), are considered discontinuous in our study (Fig. 5).
Appendix B Performance Metrics
@@
The performance metrics used in this study include Precision (Prec), Recall (Rec), True Positive Rate (TPR), False Positive Rate (FPR), F1-score, Antigen-level intersection over union (AgIoU), Matthews correlation coefficient (MCC), whose definitions are given as follows:
$\text{Prec}=\frac{\text{TP}}{\text{TP}+\text{FP}},$
$\text{Rec}=\frac{\text{TP}}{\text{TP}+\text{FN}},$
$\text{TPR}=\frac{\text{TP}}{\text{TP}+\text{FN}},$
$\text{FPR}=\frac{\text{FP}}{\text{FP}+\text{TN}},$
$\text{F1}=\frac{2\text{TP}}{2\text{TP}+\text{FP}+\text{FN}},$
$\text{AgIoU}=\frac{\text{TP}}{\text{TP}+\text{FP}+\text{FN}},$
$\text{MCC}=\frac{\text{TP}\cdot\text{TN}-\text{FP}\cdot\text{FN}}{\sqrt{(\text{TP}+\text{FP})(\text{TP}+\text{FN})(\text{TN}+\text{FP})(\text{TN}+\text{FN})}},$
@@
where TP, TN, FP and FN denote the number of true positives, true negatives, false positives and false negatives, respectively. The relationships between some of these metrics are given as:
$\text{F1}=\frac{2\cdot\text{Prec}\cdot\text{Rec}}{\text{Prec}+\text{Rec}}=\frac{2\cdot\text{AgIoU}}{1+\text{AgIoU}},$
@@
Appendix C Ablation
C.1 ViT1D
Model Embed Dim Patch Size Depth Heads MLP Ratio
ViT1D-9 1536 1 9 12 2.0
ViT1D-12 1536 1 12 12 2.0
@@ C.2 ResNet1D
input 1152 - $\checkmark$ $\checkmark$
stem 288 7 $\checkmark$ $\checkmark$
layer1 576 (144 $\times$ 4) 3 3 3
layer2 1152 (288 $\times$ 4) 3 4 8
layer3 2304 (576 $\times$ 4) 3 23 36
layer4 4608 (1152 $\times$ 4) 3 3 3
@@
We employed two variants of the ResNet1D backbone in the ablation studies: ResNet1D-101 and ResNet1D-152. As shown in Table 5, both models accept input feature maps with 1152 channels and begin with an initial convolutional layer producing 320 channels. The networks consist of four residual stages of bottlenecks with the expansion factor of 4, each stage starting by downsampling the channel dimension to half. The difference lies in the number of bottleneck blocks per stage. ResNet1D-101 contains [3, 4, 23, 3] blocks across the four layers, while ResNet1D-152 is deeper with [3, 8, 36, 3] blocks. The final output is projected to the target classes through a $\times1$ convolutional layer. .